OpenStudy (anonymous):

Match the expression in column I with those in column II with those in Column I 1. log xy a. log x + 2 log y 2. logx/y b. logx + logy/2 3. log xy^2 c. ½ log x – log y

OpenStudy (anonymous):

can anyone help me????

OpenStudy (anonymous):

do you know it @eyeh8maf

OpenStudy (anonymous):

The log rules needed to match these are the mulplication, division, and exponent of a log...do you know these rules?

OpenStudy (anonymous):

no i dont think so

OpenStudy (anonymous):

\[\log(a\cdot b) = \log(a) + \log(b) \\ \log(\frac{a}{b}) = \log(a) - \log(b) \\ \log(a^x) = x \cdot \log(a)\]